I have posted the assignment below and I know what the final numbers should be which are

CF year 0 = -84,000

CF year 1= 17,800

CF year 5 = 72,475

NPV = 10,942

I am having a really hard time with how she is coming up with these numbers

Gustav is facing a new financial decision. His old grinding and blending equipment is showing its age and requires more frequent shutdowns. He consulted with the local maintenance team, and they suggested that the equipment can be reinvigorated. At the same time, Gustav is thinking that a new machine can improve the company’s productivity, reduce maintenance costs, and decrease idle time.
Here is the information Gustav has gathered so far.
- The old grinding equipment
o It has a book value of $35,000.
o Gustavalreadyfoundabuyerforit;thesellingpriceis$35,000.
o The annual depreciation is $6,000.
o At the end of 5 years, the machine can be sold for $5,000.
o Theannualmaintenanceis$8,000peryear.
o it requires $30,000 to be spent immediately and another additional major maintenance of $6,000
in year 3.
- The new grinding equipment
o Itcosts$140,000.
o It can be depreciated over 10 years using a straight-line depreciation.
o The annual maintenance starts at $3,000 per year and increases by $ 1500 per year.
o Can be sold after 5 years for $60,000
o It will save $10,000, $10,500, $10,750, $11,000, and $11,250 in labor costs in years 1, 2, 3, 4, and
5.
o It will increase sales by $7,000 per year.
Gustav calculated his own cost of capital (discount rate), and it is 12%. The tax rate for his business is 30%. First, you must find the FCFs.
a) Using the NPV analysis, advise Gustav whether he should purchase the new machine.
b) Using the IRR analysis, advise Gustav whether he should purchase the new machine. Do you reach the
same conclusion?
c) Are there any other factors Gustav should consider when making his decision? Think about strategic,
qualitative, and other non-financial factors. Find at least 3.
d) You already know that Gustav wants a thorough investigation. Therefore, he wants to conduct these
additional analyses. Use only the NPV method for this part. What if
- The new machine can be sold for $80,000 in 5 years.
- The tax rate changes starting with year two to 40% due to new legislation. This legislation has not passed
yet.
- Sales will not increase due to a recession in the first two years, but they will increase by $4,000 starting
with year 3. Gustav believes that the economy will experience a recessionary period in the coming years.
- All of the above factors change at the same time.
e) Write a short recommendation for Gustav. Be specific on how you reached your conclusion.

Notes:
Use excel. It will save you time with the sensitivity analysis (part d).
For part d, you will have 4 scenarios: change in selling price, change in tax rate, change in sales, all changes together.
Answer #1

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The cells highlighted in yellow contain the data points you want to be answered. These figures are exactly same as what you have reported. Figures in parenthesis, if any, mean negative values. All financials are in $. Adjacent cells in blue contain the formula in excel I have used to get the final output.

I have offered explanation below for difficult line items:

  • We have calculated the cash flows under each of the two options: Old machine cash flows and New machine cash flows. We have then calculated the incremental cash flows as = New machine cash flows - Old machine cash flows.
  • We have then calculated the NPV of incremental cash flows. That's the answer for NPV.
  • Old machine sale value, Item A: The old machine can be sold currently at 35,000 against the book value of 35,000. Hence, there is no capital gain and hence there is no tax. By continuing with the old machine, we are foregoing this 35,000 amount today, to get its salvage value as $ 5,000 after five years. That's how line item A has come.
  • Line items B, C, D are straightway plotting the data from the case. E, F, G, H & I have linkage columns to explain the mathematics.
  • In case of new machine cash flows, all items are self explanatory or can be understood through the linkage column, except the line item S.
  • After five years, the old machine have a book value = Purchase cost - total depreciation on five years = 140,000 - 140,000 / 10 x 5 = 70,000. Sale value = 60,000. Hence, gain = sale value - book value = 60,000 - 70,000 = -10,000. Hence, tax on gain = 30% x gain = 30% x (-10,000) = -3,000. Hence, post tax sale value = Sale value - tax on gain = 60,000 - (- 3,000) = $ 63,000.
  • I sincerely hope your queries have been answered.

р R S T 0 1 2 3 4 5 (35,000) (6,000) (8,000) (6,000) (8,000) (6,000) (8,000) 5,000 (6,000) (8,000) (14,000) 4,200 (30,000) (3